OpenStudy (anonymous):

SOLVED THE MIN JSUT NEED HELP FIGURING OUT TO DO THE MAX. Determine the maximum value and the minimum value of f (x, y) = x2 + y2 − x − y on the closed unit disk D : x2+y2 ≤ 1. (Hint: Recall that the unit circle x2+y2 = 1 can be parametrized as x = cos t, y = sin t.)

OpenStudy (dumbcow):

when i take the derivative f'(t) = -cos(t) + sin(t) = 0 then sin(t) = cos(t) t = pi/4 is that look right?

OpenStudy (anonymous):

YES

OpenStudy (dumbcow):

then min = 1-2/sqrt(2) = -.414 and max would have to be 1 i guess

OpenStudy (dumbcow):

when t = -pi/4, sin = -cos max = 1 -(cos-cos) = 1

OpenStudy (dumbcow):

wow i was off t = 5pi/4 is also a solution to sin = cos max = 1+2/sqrt(2) = 2.414